# Hashed Watermark as a Filter: Defeating Forging and Overwriting Attacks in Weight-based Neural Network Watermarking

### v1

@@ Abstract
As valuable digital assets, deep neural networks necessitate robust ownership protection, positioning neural network watermarking (NNW) as a promising solution. Among various NNW approaches, weight-based methods are favored for their simplicity and practicality; however, they remain vulnerable to forging and overwriting attacks. To address those challenges, we propose NeuralMark, a robust method built around a hashed watermark filter. Specifically, we utilize a hash function to generate an irreversible binary watermark from a secret key, which is then used as a filter to select the model parameters for embedding. This design cleverly intertwines the embedding parameters with the hashed watermark, providing a robust defense against both forging and overwriting attacks. An average pooling is also incorporated to resist fine-tuning and pruning attacks. Furthermore, it can be seamlessly integrated into various neural network architectures, ensuring broad applicability. Theoretically, we analyze its security boundary. Empirically, we verify its effectiveness and robustness across 13 distinct Convolutional and Transformer architectures, covering five image classification tasks and one text generation task. The source codes are available at https://github.com/AIResearch-Group/NeuralMark.
Index Terms:
Neural network watermarking, weight-based approach, hashed watermark filter, neural network ownership.
I Introduction
@@
The advancements in artificial intelligence have led to the development of numerous deep neural networks, particularly large language models [1, 2, 3, 4, 5, 6]. Training such models requires substantial investments in human resources, computational power, and other resources, as exemplified by GPT-4, which costs around \$$40$ million to train [7]. Thus, they can be regarded as valuable digital assets, necessitating urgent measures for ownership protection. To this end, neural network watermarking (NNW) approaches [8, 9, 10, 11, 12] have been proposed to protect model ownership by embedding watermarks within the neural network. Methods that require access to model parameters for watermark embedding and verification fall under white-box neural network watermarking (NNW) [13, 14, 15, 16, 17, 18, 19], whereas those that do not require access to the model parameters belong to black-box NNW [20, 21, 22, 23, 24, 25, 26]. Both approaches have demonstrated significant progress in safeguarding model ownership [8] and hold promise for integration in practical applications [27, 15]. Given the distinct challenges inherent in white-box and black-box approaches, this article concentrates on white-box NNW, with black-box NNW left for future exploration.
@@
Existing white-box NNW approaches can be broadly categorized into three sub-branches: (i) Weight-based methods [13, 28, 29, 14, 18] embed watermarks into model parameters; (ii) Passport-based methods [27, 15, 16, 17] introduce passport layers to replace normalization layers for watermark embedding; and (iii) Activation-based methods [30, 31, 32] incorporate watermarks into the activation maps of intermediate layers. Among those methods, weight-based methods are appealing due to their inherent simplicity and practicality. By embedding watermarks into the model’s weights, those methods offer a straightforward process that can be seamlessly integrated into various network architectures without altering the original structure. This feature renders them valuable for various practical applications. Although several state-of-the-art weight-based methods [28, 29, 14, 18] can effectively resist fine-tuning and pruning attacks, they remain partially vulnerable to forging, overwriting, or both types of attacks.
@@
On one hand, forging attacks attempt to fabricate counterfeit watermarks and infer the corresponding secret key through reverse engineering, by freezing the model parameters. In this scenario, the adversary could claim the model’s ownership, resulting in ownership ambiguity. On the other hand, overwriting attacks aim to remove the original watermark by embedding a counterfeit one. In particular, adversaries can adaptively increase the embedding strength of their watermarks without being required to match the original watermark’s embedding strength. In such cases, the original watermark may be removed while the adversary’s watermark is embedded, leading to the invalidation of the model’s ownership. This raises a question: “How can we design a more robust and effective weighted-based method that defends against both forging and overwriting attacks?
@@
To address this challenge, we propose NeuralMark, a robust weighted-based method centered on a hashed watermark filter. Specifically, we use a hash function to generate an irreversible binary watermark from a secret key, which is then employed as a filter to select the model parameters for embedding. The avalanche effect of the hash function [33] ensures that even slight changes in the input lead to significant, unpredictable variations in the output, effectively impeding gradient calculation and making reverse engineering infeasible. Moreover, because the hashed watermarks generated by the model owner and the adversary are distinct, using them as private filters reduces the overlap in selected parameters, especially when the filtering process is performed repeatedly. This mechanism significantly increases the difficulty for adversaries to identify and manipulate the filtered parameters, thereby protecting the original watermark. Therefore, the hashed watermark filter cleverly intertwines the embedding parameters with the hashed watermark, providing a robust defense against both forging and overwriting attacks. Furthermore, we also apply an average pooling mechanism to the filtered parameters due to its resilience against fine-tuning and pruning attacks. Upon obtaining the resulting parameters, the hashed watermark is embedded into those parameters using a lightweight watermarking embedding loss, ensuring no degradation in model performance. During verification, the embedded watermark can be extracted from the model’s parameters to verify model ownership.
The main contributions of this article are highlighted below.
@@
The remainder of this article is organized as follows. We review related work in Section II. Section III introduces the preliminary, and Section IV presents the threat model. NeuralMark is detailed in Section V. Extensive experiments in Section VI evaluate its effectiveness and robustness. Finally, we make conclusions in Section VII.
II Related Work
II-A Weight-based Method
This kind of methods [13, 28, 29, 14, 18] embeds watermarks into the model parameters of neural networks. For instance, [13] proposes the first weight-based method, which embeds the watermark into the model parameters of an intermediate layer in the neural network. Another example is [29], which presents a method based on spread transform dither modulation that enhances the secrecy of the watermark. However, those two methods cannot effectively resist forging and overwriting attacks. Moreover, [28] utilizes the secret keys to pseudo-randomly select parameters for watermark embedding and apply spread-spectrum modulation to disperse the modulated watermark across different layers. This method effectively defends overwriting attacks while neglecting forging attacks. Additionally, [14] proposes to greedily choose important model parameters for watermark embedding without an additional secret key. Although this method is effective against forging attacks, it fails to provide strong resistance to overwriting attacks of varying strength levels. Recently, [18] introduces random noises into the watermarked parameters and then employs a majority voting scheme to aggregate the verification results across multiple rounds. While this method enhances the watermark’s robustness to some extent, it remains ineffective against forging and overwriting attacks.
II-B Passport-based Method
This group of methods [27, 15, 16, 17] integrates the watermark into the normalization layers in neural networks. Specifically, [27, 15] propose the first passport-based method, which utilizes additional passport samples (e.g., images) to generate affine transformation parameters for the normalization layers, tightly binding them to the model performance. Subsequently, [16] integrates a private passport-aware branch into the normalization layers, which is trained jointly with the target model and is used solely for watermark verification. Recently, [17] argues that binding the model performance is insufficient to defend against forging attacks, and thus proposes establishing a hash mapping between passport samples and watermarks.
II-C Activation-based Method
This category of methods [30, 31, 32] incorporates watermarks into the activation maps of intermediate layers in neural networks. For instance, [30] incorporates the watermark into the mean vector of activation maps generated by predetermined trigger samples. Similarly, [31] directly integrates the watermark into the activation maps associated with the trigger samples. Additionally, [32] embeds the watermark into the hidden memory state of a recurrent neural network.
Figure 1: Illustrations of different types of attacks. (a) Forging attack: the adversary aims to generate a counterfeit secret key–watermark pair without modifying the model parameters. (b) Overwriting attack: the adversary embeds a counterfeit watermark to overwrite the original one. (c) Fine-tuning attack: the adversary fine-tunes the model in an attempt to remove the original watermark. (d) Pruning attack: the adversary prunes the model parameters to remove the original watermark. In all cases, $\mathbf{P}(\mathbb{M}[\theta])$ denotes the performance of the model parameterized by $\theta$.
III Preliminary
In this section, we elaborate on the principle and vulnerability of the first weight-based method [13], which we refer to as VanillaMark, serving as the foundation for subsequent weight-based watermarking methods [28, 29, 14, 18]. Specifically, it begins by selecting, averaging, and flattening a subset of model parameters $\bm{\theta}$ into a parameter vector $\widetilde{\mathbf{w}}\in\mathbb{R}^{k}$. A secret key matrix $\mathbf{K}\in\mathbb{R}^{k\times n}$ is then used to derive the extracted watermark via $\widetilde{\mathbf{b}}=\delta(\widetilde{\mathbf{w}}\mathbf{K})$, where $\delta(\cdot)$ denotes the sigmoid function. To embed a target binary watermark $\mathbf{b}$ [^1] A watermark in this article refers to a binary vector of ones and zeros. into $\widetilde{\mathbf{w}}$, VanillaMark optimizes the following objective:
@@ III Preliminary
where $\mathcal{L}_{m}$ denotes the main task loss (e.g., classification loss), $\mathcal{L}_{e}(\cdot,\cdot)$ represents the binary cross-entropy loss, and $\lambda$ is a positive trade-off hyper-parameter. Although VanillaMark is simple and pioneering, and can resist fine-tuning and pruning attacks [13], it remains vulnerable to the following two attacks:
  • Forging Attack: An adversary can learn the secret key for any arbitrary watermark. Specifically, given a counterfeit watermark $\mathbf{b}_{a}$, the attacker can learn a corresponding key $\mathbf{K}_{a}$ by minimizing $\mathcal{L}_{e}(\mathbf{b}_{a},\widetilde{\mathbf{b}}_{a})$ while keeping the model parameters frozen, i.e., $\mathbf{K}_{a}=\arg\min_{\mathbf{K}_{a}}\mathcal{L}_{e}(\mathbf{b}_{a},\widetilde{\mathbf{b}}_{a})$.
  • Overwriting Attack: VanillaMark neither protects the confidentiality of watermarked parameters nor ensures non-overlapping usage between the model owner’s and the adversary’s parameters. Once the watermarked parameters are identified, an adversary can forge a counterfeit watermark tuple $\{\mathbf{K}_{a},\mathbf{b}_{a}\}$ and embed $\mathbf{b}_{a}$ into those parameters by optimizing the objective function: $\min_{\theta}\mathcal{L}_{m}+\lambda\mathcal{L}_{e}(\mathbf{b}_{a},\widetilde{\mathbf{b}})$. Since different watermarks often induce conflicting gradients on the same parameters, the newly embedded watermark can easily overwrite the original one.
IV Threat Model
IV-A Adversary Capabilities
We assume that an adversary can illegally obtain a watermarked model, identify the layers containing the watermark, and obtain the original training datasets, but is limited in computational resources. This constraint is reasonable, as an attacker with sufficient computational resources could simply train a new model from scratch, making model theft unnecessary. As discussed above, this work primarily focuses on forging and overwriting attacks, while also considering fine-tuning and pruning attacks. The specific threat scenarios are detailed as follows. (1) Forging Attack: As illustrated in Fig. 1a, in a forging attack, the adversary aims to generate a counterfeit secret key–watermark pair without modifying the model parameters. Specifically, the adversary first randomly forges a counterfeit watermark and then derives a corresponding secret key by optimizing it while keeping the model parameters frozen [27, 15]. (2) Overwriting Attack: As presented in Fig. 1b, in an overwriting attack, the adversary attempts to embed a counterfeit watermark to overwrite the original one [14]. (3) Fine-tuning Attack: As depicted in Fig. 1c, in a fine-tuning attack, the adversary fine-tunes the model in an attempt to remove the original watermark. (4) Pruning Attack: As shown in Fig. 1d, in a pruning attack, the adversary attempts to remove the original watermark by pruning the model parameters.
Figure 2: Illustration of hashed watermark filter. Here, the model owner’s hashed watermark is $[1,0,1,0]$, while the adversary’s is $[0,1,1,0]$. Without filtering, all 16 parameters overlap. After one round of filtering, each retains eight parameters, with four overlapping. A second round leaves four parameters each, with no overlap.
IV-B Attack Success Criteria
Building on insights from [27, 15, 34, 35], a successful attack on a watermarked model typically requires the adversary to either (i) forge a counterfeit watermark without altering the model parameters, or (ii) remove the original watermark through parameter modifications, all while preserving model performance. If the adversary only embeds a counterfeit watermark without removing the original one, the resulting model contains both. In this case, the model owner can submit a version containing only the original watermark to an authoritative third-party for verification. In contrast, the adversary cannot provide a model with only the counterfeit watermark, as the original watermark remains intact. As a result, the adversary cannot convincingly claim ownership unless they train a new model embedded solely with their own watermark. This not only makes stealing the original model unnecessary but also incurs significant training costs. Accordingly, we define the success criteria for each type of attack as follows. (1) Success Criteria for Forging Attack: Forge a counterfeit watermark that passes verification without modifying the model parameters. (2) Success Criteria for Overwriting Attack: Remove the original watermark and embed a counterfeit one by modifying the model parameters, while maintaining model performance. (3) Success Criteria for Fine-tuning Attack: Remove the original watermark through fine-tuning, while maintaining model performance. (4) Success Criteria for Pruning Attack: Remove the original watermark through parameter pruning, while maintaining model performance.
V Methodology
In this section, we present NeuralMark, a weight-based method designed to protect model ownership. The objective is to train a watermarked model $\mathbb{M}(\bm{\theta}^{*})$ on a given training dataset $\mathcal{D}$ such that the model parameters $\bm{\theta}^{*}$ embed a binary watermark $\mathbf{b}$ while satisfying the following criteria: (i) the watermark imposes negligible impact on the model performance and remains difficult for adversaries to detect; and (ii) the embedded watermark exhibits robustness against the adversarial attacks defined in Section IV.
V-A Motivation
As discussed in Section III, most weight-based methods struggle to simultaneously defend against both forging and overwriting attacks. On the one hand, forging attacks aim to generate a counterfeit watermark and derive the corresponding secret key via gradient backpropagation, while keeping the model parameters fixed. Defending against such attacks requires disrupting gradient computation to hinder reverse-engineering. On the other hand, overwriting attacks attempt to remove the original watermark by embedding a counterfeit one. Once watermarked parameters are identified, the adversary can overwrite the original watermark. Since each watermark updates the model parameters in a distinct and often conflicting direction, embedding a new watermark can easily disrupt the original one. Therefore, to resist such attacks, it is essential to keep the watermarked parameters confidential and ensure that those used by the model owner and the adversary are non-overlapping.
To address both threats, we propose a hashed watermark filter that leverages an irreversible binary watermark as a private filter to restrict embedding to a secret subset of model parameters. Specifically, we utilize a hash function to generate an irreversible binary watermark from a secret key, which is then applied as a filter to select the model parameters for embedding. This design provides two key properties:
  • Gradient Obfuscation: The avalanche effect of the hash function ensures that even minor changes in the input lead to large, unpredictable changes in the output, effectively impeding gradient computation and rendering reverse-engineering infeasible.
  • Embedding Isolation: Since the hashed watermarks derived by the model owner and the adversary are inherently distinct, using them as private filters can effectively reduce the overlap in selected parameters, especially when the filtering process is performed repeatedly. As exemplified in Fig. 2, the model owner’s hashed watermark is $[1,0,1,0]$, while the adversary’s is $[0,1,1,0]$. Without filtering, all 16 model parameters are shared, yielding a $100\%$ overlap ratio. After the first round of filtering, each party retains eight parameters, with four overlapping, reducing the overlap to $50\%$. A second filtering round results in four parameters per party, with zero overlap, achieving a $0\%$ overlap ratio. This progressive isolation ensures that as filtering continues, the overlap between the model owner’s and the adversary’s selected parameters is significantly reduced. Thus, it becomes increasingly difficult for the adversary to identify and manipulate the owner’s watermarked parameters, even when increasing the embedding strength of their watermarks, thereby preserving the integrity of the original watermark against overwriting attacks.
In summary, those properties enable the hashed watermark filter to tightly entangle the embedding parameters with the hashed watermark, providing strong resistance against both forging and overwriting attacks. This mechanism forms the core of NeuralMark, which we will elaborate on next.
V-B NeuralMark
As depicted in Fig. 3, NeuralMark consists of three primary steps: (i) hashed watermark generation; (ii) watermark embedding; and (iii) watermark verification. Next, we detail how each step works.
V-B1 Hashed Watermark Generation
As aforementioned, we construct a hash-based mapping from a secret key to a binary watermark, as shown in Fig. 3a. Formally, the watermark $\mathbf{b}\in\{0,1\}^{n}$ is generated by $\mathbf{b}=\mathcal{H}(\mathbf{K})$, where $\mathbf{K}\in\mathbb{R}^{k\times n}$ is a secret key matrix with elements drawn from a random distribution (e.g., normalized Gaussian distribution), $\mathcal{H}(\cdot)$ denotes a hash function, and $n$ indicates the length of the watermark. To accommodate various application requirements, we adopt SHAKE-256 [36], an extendable-output function from the SHA-3 family that allows dynamic adjustment of output length. Furthermore, auxiliary content $\mathcal{C}$ (e.g., textual descriptors or unique identifiers) can also be incorporated into the hash function, yielding $\mathbf{b}=\mathcal{H}(\mathbf{K}||\mathcal{C})$, where ||||| | denotes the concatenation operation. This mechanism enables context-aware watermark generation without compromising the avalanche effect of the hash function. For simplicity, we omit auxiliary content in the subsequent experiments.
V-B2 Watermark Embedding
As illustrated in Fig. 3b, to embed the hashed watermark $\mathbf{b}$ into the model $\mathbb{M}(\bm{\theta})$, we first select and flatten a subset of parameters (e.g., one-layer parameters) from $\bm{\theta}$ into a parameter vector $\mathbf{w}\in\mathbb{R}^{m}$. Then, we utilize the hashed watermark filter to select the model parameters for embedding. Specifically, let $\mathbf{w}^{(0)}=\mathbf{w}$ be the initial parameter vector. In the $r$-th ($r\in\{1,\cdots,R\}$) filtering round, the watermark $\mathbf{b}$ is repeated to match the length of $\mathbf{w}^{(r-1)}$, forming $\mathbf{b}^{(r)}$, with any excess parameters in $\mathbf{w}^{(r-1)}$ discarded. The parameter vector $\mathbf{w}^{(r)}$ is constructed by selecting the elements from $\mathbf{w}^{(r-1)}$ at positions where $\mathbf{b}^{(r)}$ equals one, i.e., $\mathbf{w}^{(r)}=\big{[}w_{i}^{(r-1)}\mid i\in\{j\mid b_{j}^{(r)}=1\}\big{]}$, where $w^{(r-1)}_{i}$ is the $i$-th element of $\mathbf{w}^{(r-1)}$, and $b^{(r)}_{j}$ is the $j$-th element of $\mathbf{b}^{(r)}$. After completing the whole watermark filtering process, the filtered parameter vector $\mathbf{w}^{(R)}$ is obtained. Next, we adopt the average pooling $\text{AVG}(\cdot)$ operation [37] to calculate the final parameters as $\widetilde{\mathbf{w}}=\text{AVG}(\mathbf{w}^{(R)})\in\mathbb{R}^{k}$. This operation aggregates parameters across broader regions, thereby enhancing robustness against parameter perturbations caused by fine-tuning and pruning attacks. Finally, we formulate the overall optimal objective as
@@ V-B2 Watermark Embedding
where $\mathcal{L}_{m}$ denotes the main task loss (e.g., classification loss), $\mathcal{L}_{e}(\cdot,\cdot)$ represents the binary cross-binary loss, $\widetilde{\mathbf{b}}=\delta(\widetilde{\mathbf{w}}\mathbf{K})$ denotes the extracted watermark, with $\delta(\cdot)$ being the sigmoid function, and $\lambda$ is a positive trade-off hyper-parameter. By minimizing Eq. (2), the watermark can be embedded into model parameters during the main task training. The watermark embedding process is summarized in Algorithm 1.
Algorithm 1 Watermark Embedding in NeuralMark
Training dataset $\mathcal{D}$, secret key $\mathbf{K}$, index of embedding layer $\mathbf{I}_{e}$, hyper-parameters $\lambda$, $T$, and filter rounds $R$.
Watermarked model $\mathbb{M}(\theta^{*})$.
Randomly initialize the model parameter $\theta$.
Generate the watermark $\mathbf{b}=\mathcal{H}(\mathbf{K})$.
for $t=0$ to $T-1$ do
     Use $\mathbf{I}_{e}$ to select a subset from $\theta$ and flatten it into $\mathbf{w}$.
     for $r=1$ to $R$ do
         Perform watermark filtering on $\mathbf{w}$ to obtain $\mathbf{w}^{(r)}$.
     end for
     Apply average pooling on $\mathbf{w}^{(R)}$ to yield $\widetilde{\mathbf{w}}$.
     Execute sigmoid mapping on $\widetilde{\mathbf{w}}\mathbf{K}$ to produce $\widetilde{\mathbf{b}}$.
     Update $\theta$ based on Eq. (2).
end for
V-B3 Watermark Verification

### v2

# Hashed Watermark as a Filter: A Unified Defense Against Forging and Overwriting Attacks in Neural Network Watermarking

@@ Abstract
As valuable digital assets, deep neural networks necessitate robust ownership protection, positioning neural network watermarking (NNW) as a promising solution. Among NNW approaches, weight-based methods embed watermarks directly into model parameters; however, they remain generally susceptible to forging and overwriting attacks. To address those challenges, we propose NeuralMark, a robust method built around a hashed watermark filter. Specifically, we utilize a hash function to generate an irreversible binary watermark from a secret key, which is then used as a filter to select the model parameters for embedding. This design cleverly intertwines the embedding parameters with the hashed watermark, providing a robust defense against both forging and overwriting attacks. Average pooling is also incorporated to resist fine-tuning and pruning attacks. Furthermore, it can be seamlessly integrated into various neural network architectures, ensuring broad applicability. We theoretically analyze its security boundary and highlight the necessity of using a hashed watermark as a filter. Empirically, we demonstrate its effectiveness and robustness across 13 distinct Convolutional and Transformer architectures, covering five image classification tasks and one text generation task. The source codes are available at https://github.com/AIResearch-Group/NeuralMark.
Introduction
@@
The advancements in artificial intelligence have led to the development of numerous deep neural networks, particularly large language models (Mann et al. 2020; Achiam et al. 2023; Bai et al. 2023; Dubey et al. 2024; Cao et al. 2024). Training such models requires substantial investments in human resources, computational power, and other resources, as exemplified by GPT-4, which costs around \$$40$ million to train (Cottier et al. 2024). Neural networks can thus be regarded as valuable digital assets, making effective ownership protection essential. Motivated by this need, neural network watermarking (NNW) methods (Li, Wang, and Barni 2021; Lukas et al. 2022; Xue et al. 2021) have been proposed. They are generally categorized into three types: (i) White-box methods require access to the model’s internal information (e.g., parameters or activations) (Uchida et al. 2017; Liu, Weng, and Zhu 2021; Fan et al. 2021; Li et al. 2024); (ii) Black-box method require querying the model’s input–output mapping (Zhang et al. 2021; Huang et al. 2023; An et al. 2025); and (iii) Box-free method require only the model outputs and are particularly suitable for image generative models (Zhang et al. 2021; Huang et al. 2023; An et al. 2025). All three categories have demonstrated significant progress in safeguarding model ownership (Sun et al. 2023; Ngo et al. 2025). Given the distinct challenges associated with each type, this work focuses on white-box NNW, leaving the investigation of other types for future work.
@@
Existing white-box NNW methods can be broadly categorized into three sub-branches: (i) Weight-based methods (Uchida et al. 2017; Feng and Zhang 2020; Li, Tondi, and Barni 2021; Liu, Weng, and Zhu 2021; Li et al. 2024) embed watermarks into model parameters; (ii) Passport-based methods (Fan, Ng, and Chan 2019; Fan et al. 2021; Zhang et al. 2020; Liu et al. 2023) introduce passport layers to replace normalization layers for watermark embedding; and (iii) Activation-based methods (Rouhani, Chen, and Koushanfar 2019; Li et al. 2021; Lim et al. 2022) incorporate watermarks into the activation maps of intermediate layers (see Appendix A for a detailed discussion of related work). Among those methods, weight-based approaches embed watermarks directly into the model’s parameters. This allows seamless integration into various network architectures without modifying the original structure (Uchida et al. 2017; Li, Wang, and Barni 2021), providing a direct and easily implementable mechanism for watermark embedding. Although several state-of-the-art weight-based methods (Feng and Zhang 2020; Li, Tondi, and Barni 2021; Liu, Weng, and Zhu 2021; Li et al. 2024) can effectively resist fine-tuning and pruning attacks, they remain partially vulnerable to forging, overwriting, or both types of attacks.
@@
On the one hand, forging attacks attempt to fabricate counterfeit watermarks and infer the corresponding secret key through reverse engineering, by freezing the model parameters. In this scenario, the adversary could claim the model’s ownership, resulting in ownership ambiguity. On the other hand, overwriting attacks aim to remove the original watermark by embedding a counterfeit one. In particular, adversaries can adaptively increase the embedding strength of their watermarks without being required to match the original watermark’s embedding strength. In such cases, the original watermark may be removed while the adversary’s watermark is embedded, leading to the invalidation of the model’s ownership. This raises a question: “How can we design a more robust and effective weighted-based method that defends against both forging and overwriting attacks?
@@
To explore this question, we propose NeuralMark, a weighted-based method centered on a hashed watermark filter. Specifically, we use a hash function to generate an irreversible binary watermark from a secret key, which is then employed as a filter to select the model parameters for embedding. The avalanche effect of the hash function (Webster and Tavares 1985) ensures that slight input changes induce significant, unpredictable output variations, impeding gradient calculation and making reverse-engineering-based forging attacks infeasible. Moreover, using distinct hashed watermarks as private filters reduces parameter overlap, especially under repeated filtering, which increases the difficulty for adversaries to locate and manipulate the embedded parameters, thereby hindering overwriting attacks. As a result, the hashed watermark filter cleverly intertwines the embedding parameters with the hashed watermark, providing a robust defense against both forging and overwriting attacks. Furthermore, we also apply an average pooling mechanism to the filtered parameters due to its resilience against fine-tuning and pruning attacks. Upon obtaining the resulting parameters, the hashed watermark is embedded into those parameters using a lightweight embedding loss. During verification, the embedded watermark is extracted to verify model ownership.
The main contributions of this paper are threefold.
@@
Threat Model
Adversary Capabilities
We assume a fully trusted third-party verifier responsible for watermark verification. An adversary can illegally access a watermarked model, identify the watermark-containing layers, and obtain the original training datasets, but is limited in computational resources. This constraint is reasonable, as an attacker with sufficient computational resources could train a new model from scratch, making model theft unnecessary. As discussed above, this work focuses on forging and overwriting attacks, while also considering fine-tuning and pruning attacks. Those threat scenarios are detailed as follows. (1) Forging Attack: The adversary aims to generate a counterfeit secret key–watermark pair without modifying the model parameters. Specifically, the adversary first randomly forges a counterfeit watermark and then derives a corresponding secret key by optimizing it while keeping the model parameters frozen (Fan, Ng, and Chan 2019; Fan et al. 2021). (2) Overwriting Attack: The adversary attempts to embed a counterfeit watermark to overwrite the original one (Liu, Weng, and Zhu 2021). (3) Fine-tuning Attack: The adversary aims to fine-tune the model to remove the original watermark. (4) Pruning Attack: The adversary attempts to remove the original watermark by parameter pruning.
Figure 1: Illustration of the hashed watermark filter. The model owner’s hashed watermark is $[1,0,1,0]$, while the adversary’s is $[0,1,1,0]$. The watermark is repeated to match the parameter length before each round of filtering. Without filtering, all 16 parameters overlap. After the first round, each watermark retains eight parameters with four overlapping; after the second round, only four parameters remain for each, with no overlap.
Attack Success Criteria
Building on insights from (Fan, Ng, and Chan 2019; Fan et al. 2021; Zhu et al. 2020; Li et al. 2022), a successful attack on a watermarked model typically requires the adversary to either (i) forge a counterfeit watermark without altering the model parameters, or (ii) remove the original watermark through parameter modifications, all while preserving model performance. If the adversary only embeds a counterfeit watermark without removing the original one, the resulting model contains both. In this case, the model owner can submit a version containing only the original watermark to an authoritative third-party for verification. In contrast, the adversary cannot provide a model with only the counterfeit watermark, as the original watermark remains intact. As a result, the adversary cannot convincingly claim ownership unless they train a new model embedded solely with their own watermark. This not only makes stealing the original model unnecessary but also incurs significant training costs. Accordingly, we define the success criteria for each type of attack as follows. (1) Success Criteria for Forging Attack: Forge a counterfeit watermark that passes verification without modifying the model parameters. (2) Success Criteria for Overwriting Attack: Remove the original watermark and embed a counterfeit one by modifying the model parameters, while maintaining model performance. (3) Success Criteria for Fine-tuning Attack: Remove the original watermark through fine-tuning, while maintaining model performance. (4) Success Criteria for Pruning Attack: Remove the original watermark through parameter pruning, while maintaining model performance.
Methodology
In this section, we present NeuralMark, a weight-based method designed to protect model ownership. The objective is to train a watermarked model $\mathbb{M}(\bm{\theta}^{*})$ on a given training dataset $\mathcal{D}$ such that the model parameters $\bm{\theta}^{*}$ embed a binary watermark $\mathbf{b}$ [^1] Watermarks in this paper are binary vectors of 0s and 1s. while satisfying the following criteria: (i) the watermark imposes negligible impact on the model performance and remains difficult for adversaries to detect; and (ii) the embedded watermark exhibits robustness against the adversarial attacks defined in the Threat Model section.
Motivation
As aforementioned, most weight-based methods struggle to defend against both forging and overwriting attacks. On the one hand, forging attacks aim to generate a counterfeit watermark and derive the corresponding secret key via gradient backpropagation, while keeping the model parameters fixed. Defending against such attacks requires disrupting gradient computation to hinder reverse-engineering. On the other hand, overwriting attacks attempt to remove the original watermark by embedding a counterfeit one. Once watermarked parameters are identified, the adversary can overwrite the original watermark. Since each watermark updates the model parameters in a distinct and often conflicting direction, embedding a new watermark can easily disrupt the original one. Defending against such attacks is essential to preserving the confidentiality of watermarked parameters and ensuring distinct parameter usage between the model owner and the adversary.
To address both attacks, we propose a hashed watermark filter, which uses an irreversible watermark generated from a secret key via a hash function as a private filter, restricting watermark embedding to a secret parameter subset. This design provides two key properties:
  • Gradient Obfuscation: The avalanche effect of the hash function ensures that even minor input changes lead to large, unpredictable output variants, effectively impeding gradient computation and rendering reverse-engineering-based forging attacks infeasible.
  • Embedding Isolation: Since the hashed watermarks of the model owner and the adversary are inherently distinct, using them as private filters can effectively reduce the overlap in selected parameters, especially when the filtering process is performed repeatedly. As exemplified in Figure 1, the model owner’s hashed watermark is $[1,0,1,0]$, while the adversary’s is $[0,1,1,0]$. Without filtering, all 16 model parameters are shared, yielding a $100\%$ overlap ratio. After the first round of filtering, each party retains eight parameters, with four overlapping, reducing the overlap to $50\%$. A second filtering round results in four parameters per party, with zero overlap, achieving a $0\%$ overlap ratio. This progressive isolation ensures that as filtering continues, the overlap between the model owner’s and the adversary’s selected parameters is significantly reduced. Thus, it becomes increasingly difficult for the adversary to identify and manipulate the owner’s watermarked parameters, even when increasing the embedding strength of their watermarks, thereby preserving the integrity of the original watermark against overwriting attacks.
In summary, these properties allow the hashed watermark filter to tightly entangle the embedding parameters with the hashed watermark, which is essential for resisting both forging and overwriting attacks (see the Security Analysis section for details). This mechanism forms the core of NeuralMark, which we will elaborate on next.
NeuralMark
NeuralMark consists of three primary steps: (i) hashed watermark generation; (ii) watermark embedding; and (iii) watermark verification. Figure 5 in Appendix C illustrates the workflow of each step. We now elaborate on each step.
Hashed Watermark Generation
As aforementioned, we construct a hash-based mapping from a secret key to a binary watermark. Formally, the watermark $\mathbf{b}\in\{0,1\}^{n}$ is generated by $\mathbf{b}=\mathcal{H}(\mathbf{K})$, where $\mathbf{K}\in\mathbb{R}^{k\times n}$ is a secret key matrix with elements drawn from a random distribution (e.g., standard Gaussian distribution), $\mathcal{H}(\cdot)$ denotes a hash function, and $n$ indicates the length of the watermark. To accommodate various application requirements, we adopt SHAKE-256 (Dworkin 2015), an extendable-output function from the SHA-3 family that allows dynamic adjustment of output length. Furthermore, auxiliary content $\mathcal{C}$ (e.g., textual descriptors or unique identifiers) can also be incorporated into the hash function, yielding $\mathbf{b}=\mathcal{H}(\mathbf{K}||\mathcal{C})$, where $||$ denotes the concatenation operation. This mechanism enables context-aware watermark generation without compromising the avalanche effect of the hash function. For simplicity, we omit auxiliary content in the experiments.
Watermark Embedding
To embed the hashed watermark $\mathbf{b}$ into the model $\mathbb{M}(\bm{\theta})$, we first select and flatten a subset of parameters (e.g., one-layer parameters) from $\bm{\theta}$ into a parameter vector $\mathbf{w}\in\mathbb{R}^{m}$. Then, we utilize the hashed watermark filter to select the model parameters for embedding. Specifically, let $\mathbf{w}^{(0)}=\mathbf{w}$ be the initial parameter vector. In the $r$-th ($r\in\{1,\cdots,R\}$) filtering round, the watermark $\mathbf{b}$ is repeated to match the length of $\mathbf{w}^{(r-1)}$, forming $\mathbf{b}^{(r)}$, with any excess parameters in $\mathbf{w}^{(r-1)}$ discarded. The parameter vector $\mathbf{w}^{(r)}$ is constructed by selecting the elements from $\mathbf{w}^{(r-1)}$ at positions where $\mathbf{b}^{(r)}$ equals one, i.e., $\mathbf{w}^{(r)}=\big[w_{i}^{(r-1)}\mid i\in\{j\mid b_{j}^{(r)}=1\}\big]$, where $w^{(r-1)}_{i}$ is the $i$-th element of $\mathbf{w}^{(r-1)}$, and $b^{(r)}_{j}$ is the $j$-th element of $\mathbf{b}^{(r)}$. After completing the whole watermark filtering process, the filtered parameter vector $\mathbf{w}^{(R)}$ is obtained. Next, we adopt the average pooling $\text{AVG}(\cdot)$ operation (Gholamalinezhad and Khosravi 2020) to calculate the final parameters as $\widetilde{\mathbf{w}}=\text{AVG}(\mathbf{w}^{(R)})\in\mathbb{R}^{k}$. This operation aggregates parameters across broader regions, thereby enhancing robustness against parameter perturbations caused by fine-tuning and pruning attacks. Finally, we formulate the overall optimal objective as
@@
where $\mathcal{L}_{m}$ denotes the main task loss (e.g., classification loss), $\mathcal{L}_{e}(\cdot,\cdot)$ represents the binary cross-entropy loss, $\widetilde{\mathbf{b}}=\delta(\widetilde{\mathbf{w}}\mathbf{K})$ denotes the extracted watermark, with $\delta(\cdot)$ being the sigmoid function, and $\lambda$ is a positive trade-off hyper-parameter. By minimizing Eq. (1), the watermark can be embedded into model parameters during the main task training. The watermark embedding process is summarized in Algorithm 1 in Appendix D.
Watermark Verification